OpenStudy (anonymous):

Evaluate the function f(x) = 3x – 2, when x = -1. A. -9 B. -5 C. 1 D. 5 Please Help Me

OpenStudy (defiance):

Okay, so we know x = -1

OpenStudy (defiance):

Rewright it. 3(-1) - 2 =

OpenStudy (anonymous):

F(x)=3(-1)-2

OpenStudy (defiance):

Order of operation multiplication comes first

OpenStudy (defiance):

3 times -1 = -3

OpenStudy (defiance):

Now subtract 2

OpenStudy (anonymous):

-3-2=-5

OpenStudy (defiance):

You get -5

OpenStudy (anonymous):

Thank you :) can you help me with a a few more?

OpenStudy (defiance):

Sure. If you'd like.

OpenStudy (anonymous):

Evaluate the function g(t) = 3/4t ,when t = 40 A.10 B.15 C.25 D.30

OpenStudy (defiance):

3/4th of 40... Is 30. Cause 1/4 of 40 is 10 10 x 3 = 30

OpenStudy (defiance):

Answer is 30.

OpenStudy (anonymous):

ok :) Evaluate the function H(x) = 16 –x, when x = -4. A.64 B.20 C.12 D-64

OpenStudy (anonymous):

I got 20

OpenStudy (defiance):

Since x = -4 Re-wright as 16 -(-4) 16 -(-4) 12

OpenStudy (defiance):

If it was multiplaction, yes it was 20. But it's subtraction

OpenStudy (anonymous):

Yes I also got 12 ;) two more then thats it :) thankk you btw

OpenStudy (defiance):

You're subtracting "-4" from 16.

OpenStudy (defiance):

Alright.

OpenStudy (defiance):

Alright, I stand correct on the 12. It is 20.

OpenStudy (anonymous):

Find k(3 for the function k(n) = 3n^2 A.36 B.24 C.10 D.49 :0

OpenStudy (defiance):

After I thought of it.

OpenStudy (anonymous):

OH ahaha okie

OpenStudy (defiance):

Cause, two -'s make a +

OpenStudy (anonymous):

So i"m correct then

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

Find k(3 for the function k(n) = 3n^2 A.36 B.24 C.10 D.49 :)

OpenStudy (defiance):

Not sure. Maybe thanksten can help?

OpenStudy (defiance):

xD.

OpenStudy (anonymous):

Ok Thanksten :D

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

ok well here is the last one :) Find w(4) for the function w(x) = 3x + 7 A.41 B.19 C.14 D.33

OpenStudy (anonymous):

I got 19

OpenStudy (defiance):

Alright, I figured it out. The funtion is w(4) x = 4. 3 x 4 = 12 12 + 7 = 19

OpenStudy (anonymous):

ok :)

OpenStudy (defiance):

I'm working on the other. but I got 27.. xD

OpenStudy (anonymous):

xD haha ok

OpenStudy (misty1212):

HI!!

OpenStudy (defiance):

Hello, misty.

OpenStudy (anonymous):

hi Misty :)

OpenStudy (misty1212):

it is "find \(k(3)\) for the function \(k(n)=3n^2\)?

OpenStudy (misty1212):

i mean is that the one you are unsure of?

OpenStudy (defiance):

Yes, thats her problem.

OpenStudy (misty1212):

ok then \[\large k(3)=3\times 3^2=27\] is correct

OpenStudy (anonymous):

27 is not a choice on there

OpenStudy (defiance):

n = 3 So, 3 to the power 2 = 9 3 times 9 = 27...

OpenStudy (misty1212):

ok maybe there is a typo in the question?

OpenStudy (defiance):

Ask your instructor.

OpenStudy (defiance):

Indeed.

OpenStudy (anonymous):

ok THANK YOU ALL SOO MUCH ;)

OpenStudy (defiance):

Yw ;)

OpenStudy (anonymous):

Yh I also got 27